# Silence Speaks Volumes: A New Paradigm for Covert Communication via History Timing Patterns

### v1

@@ 6 Discussion
The most fundamental limitation of SHP is that its bandwidth and accuracy depend on the amount and timing of overt traffic which can be observed by communicating agents. This limitation can be alleviated by employing adaptive algorithms that dynamically adjust the bitlength, rehashing and ECC strategies based on observed network traffic conditions. Implementing SHP closer to the hardware, for example, using kernel access, could further improve performance by minimizing delays. For environments that do have a lot of overt traffic, a possible improvement is channel multiplexing: dividing overt traffic into multiple parallel channels increases spacing per channel, thereby deconflicting pointers. These multiple channels could transmit covert traffic in parallel or for different applications, leading to a substantial increase in overall bandwidth. Optimization of such a multi-channel method could essentially follow Wendzel and Keller [wendzelLowAttention].
@@
While SHP (like all other covert channels) can potentially be misused, we believe its advancement requires public in-depth studies for two reasons: (i) History covert channels have already shown to be applicable for censorship circumvention (see OPPRESSION [AsiaCCS24:OPPR]), which aids the exchange of critical opinions through networks and, for this reason, supports an essential human right. (ii) When research results on network covert channels are made public, they aid the whole community, not just some of the actors (e.g., censors), allowing everybody to develop protection mechanisms. In contrast, when censors or cyber criminals develop such novel methods, they do not share their insights with the public, resulting in a disbalance of available knowledge.
@@ 7 Conclusion
Declarations
CRediT authorship contribution statement
Christoph Weissenborn: Conceptualization, Methodology, Software, Validation, Formal analysis, Investigation, Data curation, Visualization, Writing – original draft. Steffen Wendzel: Supervision, Project administration, Conceptualization, Methodology, Validation, Writing – review & editing.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data and code availability
All code and experimental data supporting the findings of this study are available for reproducibility and further research: https://github.com/gXeeXqBHuHDFTaEnff3Z/SHP.
Funding
This research received no external funding.
@@ Acknowledgments
We thank Tobias Schmidbauer, Sebastian Zillien, and Jörg Keller for providing code and data of DYST [DYST22], which we used for comparison. We also thank Haozhi Li, Tian Song and Yating Yang for their work on generic and sensitive anomaly detection [GAS22], which we used as approach for out ML-detection model.

### v2

@@ 6 Discussion
The most fundamental limitation of SHP is that its bandwidth and accuracy depend on the amount and timing of overt traffic which can be observed by communicating agents. This limitation can be alleviated by employing adaptive algorithms that dynamically adjust the bitlength, rehashing and ECC strategies based on observed network traffic conditions. Implementing SHP closer to the hardware, for example, using kernel access, could further improve performance by minimizing delays. For environments that do have a lot of overt traffic, a possible improvement is channel multiplexing: dividing overt traffic into multiple parallel channels increases spacing per channel, thereby deconflicting pointers. These multiple channels could transmit covert traffic in parallel or for different applications, leading to a substantial increase in overall bandwidth. Optimization of such a multi-channel method could essentially follow Wendzel and Keller [61].
@@
While SHP (like all other covert channels) can potentially be misused, we believe its advancement requires public in-depth studies for two reasons: (i) History covert channels have already shown to be applicable for censorship circumvention (see OPPRESSION [72]), which aids the exchange of critical opinions through networks and, for this reason, supports an essential human right. (ii) When research results on network covert channels are made public, they aid the whole community, not just some of the actors (e.g., censors), allowing everybody to develop protection mechanisms. In contrast, when censors or cyber criminals develop such novel methods, they do not share their insights with the public, resulting in a disbalance of available knowledge.
@@ Acknowledgments
We thank Tobias Schmidbauer, Sebastian Zillien, and Jörg Keller for providing code and data of DYST [62], which we used for comparison. We also thank Haozhi Li, Tian Song and Yating Yang for their work on generic and sensitive anomaly detection [35], which we used as approach for out ML-detection model.